The length of a rectangle exceeds its width by 8 inches. The perimeter is 80inches.find the area of the rectangle.
Answer by TimothyLamb(4379)

---
x = length
y = width
---
x = y + 8
2x + 2y = 80
---
put the system of linear equations into standard form
---
x - y = 8
2x + 2y = 80
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = length = 24 in
y = width = 16 in
a = 16*24 = 384 sq.in
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php